OpenStudy (anonymous):

Find the exact value by using a half-angle identity. cos(5π/12) I know someone knows how to do this! if you could explain too, that'd be great!

OpenStudy (anonymous):

os 5π/12 cos ((5 pi/6)/2) = +/- sqrt( (1+cos (5 pi/6))/2 ) cos(5 pi/6) = -sqrt(3)/2 cos (5 pi/12) = +/- sqrt( (2/2-sqrt(3)/2)/2 ) cos (5 pi/12) = +/- sqrt( ((2-sqrt(3))/2)/2 ) cos (5 pi/12) = sqrt( ( (2-sqrt(3))/4 ) ) 5 pi/12 is in quadrant 1 so we choose the + sign because cos is positive there

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

wait did you like re-expand the problem after cos(5 pi/6) = -sqrt(3)/2

OpenStudy (anonymous):

most of these problems are some identities written on the unit circle.

OpenStudy (anonymous):

okay.. well can you please explain each of the steps this person just gave me? i would reaaaalllyyyy appreciate it

OpenStudy (anonymous):

ok so 5pi/12 is 75 degrees rite

OpenStudy (anonymous):

i guess.. lol is it?

OpenStudy (anonymous):

pi is 180 degrees.lol you know that rite and u muliply tht by 5 and get 900 and divide 12 and you get 75 degrees

OpenStudy (anonymous):

okay but what does the 75 degrees have to do with the formula?

OpenStudy (anonymous):

just to make ur question simpler... for example lets say it asked u for the cos of pi/6 which is 30 degrees.

OpenStudy (anonymous):

there is some thing called the unit circle and it has this angles from 30,45,60,90 upto 360.. and since we have cos of Pi/6 which is 30 degrees. we can look at the unit circle for the cosine of 30 degrees and find it

OpenStudy (anonymous):

@Dwade03 WHATS THE ANSWER?! D:

OpenStudy (anonymous):

how is it done using half angle identity formula?!

OpenStudy (anonymous):

ok so since they asked for cos(75) we can use the formula of using two angles to create that angle of 75. in this case we can use cos(45+35). u get tht?

OpenStudy (lukecrayonz):

@ganeshie8

OpenStudy (anonymous):

so what in the world is the final answer? Is it cos(75)?